OpenStudy (yellowlegoguy99):

Match each actual value with the predicted value that results in a thirty−five percent error. Terms Actual: 140 Actual: 100 Actual: 80 Actual:120 Definitions Predicted: 91 Predicted: 78 Predicted: 65 Predicted: 108

OpenStudy (yellowlegoguy99):

@terenzreignz

OpenStudy (yellowlegoguy99):

@Foreverisabell

OpenStudy (yellowlegoguy99):

Do you understand the question terenz?

OpenStudy (yellowlegoguy99):

@skullpatrol Do you understand it?

OpenStudy (skullpatrol):

Understand what?

OpenStudy (yellowlegoguy99):

The question

OpenStudy (skullpatrol):

I know how to find 35% of a number.

OpenStudy (yellowlegoguy99):

Me too but what number do I have to find 35% of?

OpenStudy (skullpatrol):

The actual

OpenStudy (yellowlegoguy99):

35% of 140 = 49 35% of 100 = 35 35% of 80 = 28 35% of 120 = 42 But those aren't choices!?

OpenStudy (yellowlegoguy99):

@ns_5

OpenStudy (skullpatrol):

Yes

OpenStudy (yellowlegoguy99):

So how do I find the answer...

OpenStudy (skullpatrol):

Let's take 35% of 100 = 35...so we need to find which predicted value results in a thirty−five percent error.

OpenStudy (skullpatrol):

100 + 35 0r 100 - 35

OpenStudy (yellowlegoguy99):

100 - 35 = 65

OpenStudy (skullpatrol):

Yes, that is on the list, right?

OpenStudy (yellowlegoguy99):

Yeah

OpenStudy (skullpatrol):

Can you do the rest yourself now?

OpenStudy (yellowlegoguy99):

I don't understand the rest though...

OpenStudy (skullpatrol):

OK take 35% of 140 = 49

OpenStudy (skullpatrol):

140 + 49 Or 140 -49

OpenStudy (yellowlegoguy99):

Ohhh 140 - 49

OpenStudy (skullpatrol):

Yes 91 is there, right?

OpenStudy (yellowlegoguy99):

And the rest: 35% of 80 = 28. 80 + 28 = 108 35% of 120 = 42. 120 - 42 = 78

OpenStudy (yellowlegoguy99):

Thanks dude

OpenStudy (skullpatrol):

Thanks for trying to understand dude :)